When light from a red laser, with a wavelength of 632 nm, is incident on a certain double slit, a particular interference pattern is observed. The laser light is replaced by a beam of electrons, all with the same energy, and exactly the same interference pattern is observed. What is (a) the wavelength, (b) frequency, and (c) energy of the electrons?

1
Expert's answer
2022-04-18T15:18:54-0400

a)

"\lambda=632~nm,"

b)

"\nu=\frac c\lambda=4.7\cdot 10^{14}~Hz,"

c)

"E=h\nu=3.1\cdot 10^{-19}~J."